OpenStudy (anonymous):

Need help to factorize 2m^12-1/8?

OpenStudy (anonymous):

what do you think is common?

OpenStudy (anonymous):

i suppose the 2 is a common factor

OpenStudy (anonymous):

factoring means leaving something in the simplest term think of something else that can be factored?

OpenStudy (anonymous):

2m^12-1/8 = 2 ( m^12 - 1/16) note that m^12 - 1/16 is of the form x^12 - y^2 = (x^6 y )(x^6 - y) then you get: 2 ( m^12 - 1/16) = 2 (m^6 1/4) (m^6 - 1/4) = 2 (m^6 1/4) (m^3 1/2) (m^3 - 1/2) this is what i got, but am unsure about it

OpenStudy (anonymous):

\[\frac{ 1 }{ 8} ( m ^{12} - 1 )\]

OpenStudy (anonymous):

sorry, it's supposed to be 16m^12 - 1

OpenStudy (anonymous):

im abit confuse

OpenStudy (anonymous):

i dont really know how to explain it and i cant show yu cuz of the OS rules

OpenStudy (anonymous):

OS rules?

OpenStudy (anonymous):

open study

OpenStudy (anonymous):

okay thank you